OpenStudy (anonymous):

Express answer in exact form. An equilateral triangle is inscribed in a circle with a radius of 6". Find the area of the segment cut off by one side of the triangle. HELP ME!!!

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

\[12 \pi - 9 \sqrt {3}\]

OpenStudy (anonymous):

Give me a minute to get the explanation together.

OpenStudy (anonymous):

The area of the arc of the circle is 1/3 of the circle which is: \[(1/3) \pi (6)^{2} = 12 \pi\] The area of the triangle is a bit trickier. If you draw the triangle you can divide it into 3 smaller triangles using the radius. These are 30-120-30 triangles. We're going to break them in two to get the area exactly. So we have 2 triangles or 2*(1/2)Base*height. The hypotenuse is 6 in each case. We need to calculate the opposite: \[sin(30) = x/6 = 1/2\] \[6(1/2) = x = 3\] Now to calculate the adjacent: \[cos(30) = x/6 = \sqrt {3} /2\] \[6 \sqrt {3} / 2 = x = 3 \sqrt {3}\] The area of the 2 smaller triangles is: \[3*3 \sqrt {3} = 9 \sqrt {3} \] So we now subtract this from the arc area: \[12 \pi - 9 \sqrt {3} \]

OpenStudy (anonymous):

Keep in mind that \[sin(30) = 1/2\] is a unit circle value. \[cos(30) = \sqrt {3}/2\] is also a unit circle value. You have a triangle with a hypotenuse of 2, adjacent of sqrt 3 and an opposite of 1.